OpenStudy (anonymous):

Give the coordinates of the ff. Trigonometric points. 1.) P (25pi/2) Please help me and explain how to get the coordinate of this...

OpenStudy (anonymous):

is it \(\frac{25\pi}{2}\) ?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

coterminal with \(\frac{\pi}{2}\) so just use that one

OpenStudy (anonymous):



OpenStudy (anonymous):

no you are being asked for the coordinates of that point it it the unit circle

OpenStudy (anonymous):

lets back up a second do you have any idea how to got that \[\frac{25\pi}{2}\] is "conterminal" with \[\frac{\pi}{2}\]?

OpenStudy (anonymous):

if the answer is "no" that is fine, i can explain it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok lets go slowly

OpenStudy (anonymous):

the entire circumference of the unit circle (circle with center \((0,0)\) and radius \(1\) is \[C=2\pi\]

OpenStudy (anonymous):

once you traverse the circle \(2\pi\) units, you start over again

OpenStudy (anonymous):